how much current is needed to plate 12g of Iron metal on the cathode using FeSO4 Solution in 4 hours?
1
Expert's answer
2020-01-06T04:50:44-0500

The oxidation state of iron in FeSO4 is +2

hence equivalent weight of Fe ="M.W.\/2"

where M.W. is the molecular weight = 55.8/2 = 27.9 grams

For deposition of one equivalent of metal one Farad charge is needed

For deposition of 12 grams of iron charge needed =(12/27.9 )F

converting into coulombs= (12/27.9)*96500 = 41505 C

time given = 4 hours

converting into seconds

time = 4*60*60 = 14400 seconds

Current = charge / time

Current = 41505/14400 = 2.88 Amperes